# Small time asymptotics of spectral heat content of isotropic processes

### v1

@@ Abstract.
We provide a general approach for proving small time asymptotic of spectral heat content for any translation invariant isotropic process satisfying negligible tail probability condition. As a consequence, we recover several existing results in the context of Lévy processes and Gaussian processes, and provide spectral heat content asymptotic for a class of time-changed Brownian motions.
@@
spectral heat content; isotropic processes; asymptotic behavior
@@ 1. Introduction and main result
Let $X=(X_{t})_{t\geqslant 0}$ be a stochastic process on $\mathbb{R}^{d}$ with càdlàg paths and $\mathbb{P}_{x}$ denotes the probability measure when $X_{0}=x$. For any bounded open set $\Omega\subset\mathbb{R}^{d}$, the spectral heat content of $X$ in the domain $\Omega$ is defined as
@@
$\lim_{t\to 0}Q^{X}_{\Omega}(t)=\mathrm{Vol}(\Omega),$
@@
$\lim_{t\to 0}\mu(t)=0.$
Throughout the article, we assume the negligible tail probability condition.
@@
$\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)=0.$
The goal of this article is to prove the following asymptotic of the spectral heat content of $X$.
@@
Let $\Omega$ be bounded open subset of $\mathbb{R}^{d}$ with $C^{1,1}$-boundary, and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic stochastic process satisfying Assumption 1.1. Then,
(1.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\mathrm{Per}(\Omega),$
@@
The $C^{1,1}$ regularity condition on the boundary of $\Omega$ inherits from the differentiability of the signed distance function from the boundary, see Remark 6.6 for details.
Corollary 1.3.
Let $\Omega$ be a bounded, open set with $C^{1,1}$-boundary and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic process such that
$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant T}X^{(1)}_{s}\right)^{\alpha}\right]<\infty$
for some $T>0$ and $\alpha>1$, and
$\lim_{t\to 0}\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]}=0.$
@@
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega),$
where $m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right$].
When $X=B$, the standard Brownian motion on $\mathbb{R}^{d}$, the spectral heat content can be realized in terms of the Dirichlet heat equation. More precisely, if $u(t,x)$ solves the PDE
$\frac{\partial u}{\partial t}u(x,t)$ $=\frac{1}{2}\Delta u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
@@
then the spectral heat content on $\Omega$ for $B$ is given by
$Q^{B}_{\Omega}(t)=\int_{\Omega}u(x,t)dx.$
@@
Small time asymptotic of $Q^{B}_{\Omega}(t)$ was first proved by Berg, see [12, Theorem 6.2], for domains with $R$-smooth boundary, and subsequently, a second order expansion of $Q^{B}_{\Omega}(t)$ near zero was proved in [14, Theorem 1.2] for domains with $C^{3}$-boundary. We also refer to [13] for small time asymptotic of $Q^{B}_{\Omega}$ when $B$ is the Brownian motion on a Riemannian manifold and $\Omega$ has smooth boundary. The spectral heat content asymptotic for symmetric $\alpha$-stable was obtained in [9], and it was later generalized to isotropic Lévy processes in [4], and to symmetric Lévy processes in a recent article [6]. We also refer to [5] for the spectral heat content asymptotic of Gaussian processes in $\mathbb{R}^{d}$.
In the present article, we provide a general approach for spectral heat content asymptotic that works for any translation invariant isotropic process satisfying Assumption 1.1. This assumption holds true for any isotropic Lévy process with regularly varying Lévy-Khintchine exponent, and Gaussian processes. Moreover, in Section 2.3, we provide examples of isotropic processes which are neither Markov nor Gaussian and Theorem 1.1 holds for such processes.
@@
Our method also extends to bounded domains with boundary being level sets of a $C^{1,\kappa}$ functions for $0<\kappa<1$. In all the aforementioned references [12, 13, 14, 9, 4, 5], the assumption of $C^{1,1}$-regularity is crucial as the proofs rely heavily on the uniform interior and exterior ball condition of the domain. In fact, from [7, Corollary 2], the uniform interior and exterior ball condition is equivalent to the $C^{1,1}$ smoothness of the boundary. For $C^{1,\kappa}$-boundary with $0<\kappa<1$, the uniform ball condition does not hold. Theorem 6.1 provides upper and lower bound of the spectral heat content for domains with boundaries given by level sets of $C^{1,\kappa}$ functions with $0<\kappa<1$. The question of validity of (1.2) for such domains still remains open.
We propose a novel approach by considering a functional version of spectral heat content. For any nonnegative integrable function $f:\mathbb{R}^{d}\to\mathbb{R}$ we define
(1.3) $Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx,$
@@
In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$ for all $t\geqslant 0$. First, we obtain asymptotic formula for $Q^{X}_{f}$ when $f$ is a compactly supported smooth function, and the lower bound of the limit in (1.2) is proved using an approximation argument. It turns out that the lower bound of (1.2) holds for any bounded, open subset with finite perimeter. The main idea behind proving the upper bound of the limit in (1.2) stems from the observation that any bounded domain with $C^{1,1}$-boundary can be expressed as $\phi^{-1}(0,\infty)$ for some bounded $C^{1,1}$ function $\phi$ with $|\nabla\phi|$ being constant on the boundary.
@@
The rest of the paper is organized as follows: in Section 2 we provide examples including isotropic Lévy processes, Gaussian processes, and a class of time-changed Brownian motions for which spectral heat content asymptotic follow directly from our main result. Section 3 provides some preliminaries about functions of bounded variation and regularity of boundary. In Section 4, we prove the asymptotic of the functional of the spectral heat content defined in (1.3). Finally, Theorem 1.1 is proved in Section 6.
@@ 2.1. Isotropic Lévy processes
We consider the class of Lévy processes studied in [4]. Let $X=(X_{t})_{t\geqslant 0}$ be an isotropic Lévy process with Lévy-Khintchine exponent $\Psi(\xi)=\psi(|\xi|)$ for any $\xi\in\mathbb{R}^{d}$, and assume that $\psi:[0,\infty)\to[0,\infty)$ is regularly varying with exponent $\beta\in[1,2]$. Then, by [10, Equation (3.2)], for any $\varepsilon>0$,
$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant c_{\varepsilon}t$
for some constant $c_{\varepsilon}>0$. On the other hand, using the bounds in [6, Equation (3.8)], we have $\lim_{t\to 0}\frac{t}{\mu(t)}=0$, which shows that Assumption 1.1 holds. Therefore, Theorem 1.1 holds for $X$, and we recover [6, Theorem 2.1(1)] for isotropic Lévy processes. In particular, when $X_{t}$ is the symmetric $\alpha$-stable process, as $t\to 0$ we have
$\mu(t)\sim\begin{cases}t^{1/\alpha}\mathbb{E}[\sup_{0\leqslant s\leqslant 1}X^{(1)}_{s}]&\mbox{if $1<\alpha\leqslant 2$}\\ \frac{1}{\pi}t\log(1/t)&\mbox{if $\alpha=1$}.\end{cases}$
For $\alpha\in(1,2]$, the asymptotic behavior of $\mu(t)$ above follows directly from the self-similarity of $\alpha$-stable process, and for $\alpha=1$, we refer to [2, Proposition 4.3(i)]. This recovers [9, Theorem 1.1] when $\alpha\geqslant 1$.
@@ 2.2. Gaussian processes
$m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right],\quad\sigma^{2}(t)=\sup_{0\leqslant s\leqslant t}\mathrm{Var}\left[(X^{(1)}_{s})^{2}\right].$
It is known that $m(t)<\infty$ for all $t\geqslant 0$. By Borell-TIS inequality, see [3, Theorem 2.1.1], for any $x>0$
(2.1) $\mathbb{P}_{0}\left(\left|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}-m(t)\right|>x\right)$ $\leqslant 2e^{-x^{2}/2\sigma^{2}(t)},$
@@
$\mathrm{Var}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]\leqslant\int_{0}^{\infty}4xe^{-x^{2}/2\sigma^{2}(t)}=4\sigma^{2}(t).$
Therefore, $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{2}\right]\leqslant m(t)^{2}+4\sigma^{2}(t)$. From [4, Equation (29)] it is known that $\lim_{t\to 0}{\sigma^{2}(t)}/{m(t)}=0$. Applying Corollary 1.3, we obtain the following result.
@@
If $\Omega$ is a bounded open set with $C^{1,1}$-boundary then,
(2.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega).$
@@
In [4, Theorem 4.4] under the condition that the rescaled process $Y^{(t)}_{u}=\mu^{-1}_{t}X^{(1)}_{tu}$ converges weakly to a Gaussian process $(Y_{u})_{u\geqslant 0}$ as $t\downarrow 0$ in Skorokhod $J_{1}$ topology, it is proved that
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathbb{E}\left[\sup_{0\leqslant u\leqslant 1}Y_{u}\right]\mathrm{Per}(\Omega),$
with $\mathbb{E}[\sup_{0\leqslant u\leqslant 1}Y_{u}]\leqslant 1$. Proposition 2.1 strengthens [4, Theorem 4.4].
2.3. Time changed Brownian motion
Consider the $d$-dimensional standard Brownian motion $B=(B_{t})_{t\geqslant 0}$ and a nonnegative, increasing continuous process $U=(U_{t})_{t\geqslant 0}$ which is independent of $B$ and $U_{0}=0$. Define $X_{t}=B_{U_{t}}$. Then, $X=(X_{t})_{t\geqslant 0}$ is translation invariant and isotropic. Let us assume that for some $\alpha>1/2$,
(2.3) $\mathbb{E}[U^{\alpha}_{t}]<\infty\quad\forall t\geqslant 0,\quad\lim_{t\to 0}\frac{\mathbb{E}[U^{\alpha}_{t}]}{\mathbb{E}[\sqrt{U_{t}}]}=0.$

### v3

@@ Abstract.
The spectral heat content of a domain $\Omega\subset\mathbb{R}^{d}$ corresponding to a $d$-dimensional stochastic process $X=(X_{t})_{t\geqslant 0}$ is defined as
$Q^{X}_{\Omega}(t)=\int_{\mathbb{R}^{d}}\mathbb{P}_{x}(\tau^{X}_{\Omega}>t)dx,$
where $\tau^{X}_{\Omega}$ is the first exit time of $X$ from $\Omega$. We provide a novel technique for proving small time asymptotic of spectral heat content for any translation invariant isotropic process satisfying negligible tail probability condition. As a consequence, we recover several existing results in the context of Lévy processes and Gaussian processes, and provide spectral heat content asymptotics for a class of $\alpha$-stable Lévy processes time-changed by right inverse of positive, increasing, self-similar Markov processes. The latter has connection to some Cauchy problems that are non-local in both time and space.
@@
spectral heat content; isotropic processes; self-similarity; asymptotic behavior
@@ 1. Introduction and main result
1.1. Main result
Let $X=(X_{t})_{t\geqslant 0}$ be a stochastic process on $\mathbb{R}^{d}$ with càdlàg paths, and let $\mathbb{P}_{x}$ denotes the probability measure when $X_{0}=x$. For any bounded open set $\Omega\subset\mathbb{R}^{d}$, the spectral heat content (SHC) of $X$ in the domain $\Omega$ is defined as
@@ 1.1. Main result
$\lim_{t\downarrow 0}Q^{X}_{\Omega}(t)=\mathrm{Vol}(\Omega),$
@@
$\lim_{t\downarrow 0}\mu(t)=0.$
Throughout the article, we assume the negligibility of the tail probability in the following sense.
@@
$\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)=0.$
The aim of this paper is to prove the following asymptotic behavior of the spectral heat content of $X$.
@@
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{d}$ with $C^{1,1}$ boundary, and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic stochastic process satisfying Assumption 1.1. Then,
(1.2) $\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\mathrm{Per}(\Omega),$
@@
The $C^{1,1}$ regularity condition on the boundary of $\Omega$ inherits from the differentiability of the signed distance function from the boundary, which is crucial in our proof of the above theorem. We refer to Remark 6.6 for details.
Assumption 1.1 is satisfied by a wide range of isotropic processes. The next result provides sufficient conditions for Assumption 1.1 to hold in terms of the moments of the supremum process $\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}$. For the ease of writing, we introduce the following notation to be used throughout the rest.
Notation 1.1.
For any stochastic process $Z=(Z_{t})_{t\geqslant 0}$ on $\mathbb{R}$, we denote the running supremum of $Z$ by $\overline{Z}$, that is,
$\overline{Z}_{t}=\sup\{Z_{s}:0\leqslant s\leqslant t\}.$
Proposition 1.3.
Let $\Omega$ be a bounded, open set with $C^{1,1}$ boundary and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic process such that
$\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{T}\right)^{p}\right]<\infty$
for some $T>0$ and $p>1$, and
$\lim_{t\downarrow 0}\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{p}\right]}{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right]}=0.$
@@
$\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right]}=\mathrm{Per}(\Omega).$
1.2. Previous wokrs
The spectral heat content in the classical sense is defined via heat equation with Dirichlet boundary condition on $\Omega$. More precisely, if $u(t,x)$ solves the PDE
$\frac{\partial}{\partial t}u(x,t)$ $=\Delta_{x}u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
@@ 1.2. Previous wokrs
then the spectral heat content of $\Omega$ is given by
$Q_{\Omega}(t)=\int_{\Omega}u(x,t)dx.$
@@
From standard theory of Markov processes, one can write $Q_{\Omega}(t)=Q^{B}_{\Omega}(t)$, where $B$ is the standard Brownian motion on $\mathbb{R}^{d}$. Small time asymptotics of $Q^{B}_{\Omega}(t)$ was first proved by van den Berg and Davies, see [25, Theorem 6.2], for domains with $R$-smooth boundary, and subsequently, van den Berg and LeGall [27, Theorem 1.2] obtained a second order expansion of $Q^{B}_{\Omega}(t)$ for domains with $C^{3}$ boundary as follows:
$Q^{B}_{\Omega}(t)=\mathrm{Vol}(\Omega)-\sqrt{\frac{2t}{\pi}}\sigma(\partial\Omega)+\frac{t}{4}\int_{\partial\Omega}H_{\Omega}d\sigma+\mathrm{O}(t^{3/2}),$
where $\sigma$ is the surface measure on $\partial\Omega$, and $H_{\Omega}$ is the mean curvature at the boundary. The above formula was further generalized to Riemannian manifolds by van den Berg and Gilkey [26] for domains with smooth boundary. Using PDE techniques, Savo [23] proved an asymptotic expansion of $Q_{\Omega}(t)$ of arbitrary order for any domain with smooth boundary in a Riemannian manifold.
While the asymptotics for Brownian motion is known for several years, results for jump processes have been obtained quite recently. In the one-dimensional case, SHC asymptotic for symmetric $\alpha$-stable process was obtained by Valverde [2], and later generalized to isotropic $\alpha$-stable processes on $\mathbb{R}^{d}$ by Park and Song [18]. We refer to [10] for SHC of isotropic Lévy processes with regularly varying Lévy-Khintchine exponents, and to a recent article [13] for symmetric Lévy processes (may not be isotropic) with some conditions on the Lévy-Khintchine exponent. When the Lévy process has bounded variation, the small time asymptotic of SHC was proved in [7], and in this case, one recovers the fractional perimeter in the limit. Beyond Lévy processes, we refer to [11, 9, 8] for separate treatments of SHC asymptotics of time-changed killed Brownian motion, time-changed killed $\alpha$-stable processes, and Gaussian processes in $\mathbb{R}^{d}$.
1.3. Our contribution
In the present article, we propose a novel approach based on simple facts from geometric measure theory for proving SHC asymptotic for any translation invariant isotropic process satisfying Assumption 1.1. This assumption holds true for any isotropic Lévy process with regularly varying Lévy-Khintchine exponent, and Gaussian processes. Moreover, in Section 2.3, we provide a wide range of examples of time-changed isotropic $\alpha$-stable processes that satisfy Assumption 1.1. As a consequence, we recover the results in [18, 8, 9, 10], and in Section 2.3.2, we provide examples where the SHC is related to some integro-partial differential equations that are non-local in both time and space. Another important aspect of our method is that it requires very minimal assumption on the isotropic process in terms of the tail probability behavior of the supremum of the one dimensional marginals. This observation led us to adapt this technique for heat content of fractional sub-Laplacian on Carnot groups and we refer the interested reader to [21].
@@ 1.3. Our contribution
Our techniques can be applied to bounded domains with boundary being level sets of a $C^{1,\kappa}$ functions for $0<\kappa<1$. In all the aforementioned references [25, 26, 27, 18, 10, 11], the assumption of $C^{1,1}$ regularity is crucial as the proofs rely heavily on the uniform interior and exterior ball condition of the domain. In fact, from [14, Corollary 2], the uniform interior and exterior ball condition is equivalent to the $C^{1,1}$ smoothness of the boundary. For $C^{1,\kappa}$ boundary with $0<\kappa<1$, the uniform ball condition does not hold. In Theorem 6.1 we provide upper and lower bound of the SHC for domains with boundaries given by level sets of $C^{1,\kappa}$ functions with $0<\kappa<1$. The question of validity of (1.2) for such domains still remains open.
Let us now briefly describe the main idea behind our approach to prove Theorem 1.1. For any nonnegative integrable function $f:\mathbb{R}^{d}\to\mathbb{R}$ we define
(1.3) $Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
@@
In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$ for all $t\geqslant 0$. Using Taylor expansion of $f$, we first obtain the asymptotic of $Q^{X}_{f}$ when $f$ is a compactly supported smooth function, and the lower bound of the limit in (1.2) is proved using an approximation argument. It turns out that the lower bound of (1.2) holds for any bounded, open subset with finite perimeter, see Proposition 5.1. The key idea to derive the upper bound of the limit in (1.2) stems from the observation that any bounded domain with $C^{1,1}$ boundary can be expressed as $\phi^{-1}(0,\infty)$ for some bounded $C^{1,1}$ function $\phi$ with $|\nabla\phi|$ being constant on the boundary.
@@
The rest of the paper is organized as follows: in Section 2 we provide examples including isotropic Lévy processes, Gaussian processes, and a class of time-changed $\alpha$-stable isotropic Lévy processes for which spectral heat content asymptotics follow directly from our main result. Section 3 provides some preliminaries about functions of bounded variation and regularity of boundary. In Section 4, we prove the asymptotic of the functional of the spectral heat content defined in (1.3). Finally, Theorem 1.1 is proved in Section 6.
@@ 2.1. Isotropic Lévy processes
We consider the class of Lévy processes studied in [10]. Let $X=(X_{t})_{t\geqslant 0}$ be an isotropic Lévy process with Lévy-Khintchine exponent $\Psi(\xi)=\psi(|\xi|)$ for any $\xi\in\mathbb{R}^{d}$, and assume that $\psi:[0,\infty)\to[0,\infty)$ is regularly varying with exponent $\beta\in[1,2]$. Then, by [20, Equation (3.2)], for any $\varepsilon>0$,
$\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\right)\leqslant c_{\varepsilon}t$
for some constant $c_{\varepsilon}>0$. On the other hand, using the bounds in [13, Equation (3.8)], we have $\lim_{t\downarrow 0}\frac{t}{\mu(t)}=0$, which shows that Assumption 1.1 holds. Therefore, Theorem 1.1 holds for $X$, and we recover [13, Theorem 2.1(1)] for isotropic Lévy processes. In particular, when $X$ is the isotropic $\alpha$-stable Lévy process, that is, $\Psi(\xi)=|\xi|^{\alpha}$, as $t\downarrow 0$ we have
$\mu(t)\sim\begin{cases}t^{1/\alpha}\mathbb{E}[\overline{X}^{(1)}_{t}]&\mbox{if $1<\alpha\leqslant 2$},\\ \frac{1}{\pi}t\log(1/t)&\mbox{if $\alpha=1$}.\end{cases}$
For $\alpha\in(1,2]$, the asymptotic behavior of $\mu(t)$ above follows directly from the self-similarity of $\alpha$-stable process, and for $\alpha=1$, we refer to [2, Proposition 4.3(i)]. This recovers [18, Theorem 1.1] when $\alpha\geqslant 1$.
@@ 2.2. Gaussian processes
$m(t)=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right],\quad\sigma(t)^{2}=\sup_{0\leqslant s\leqslant t}\mathrm{Var}\left[(X^{(1)}_{s})^{2}\right].$
We assume that $m(t)<\infty$ for all $t\in[0,1]$, which is known to be equivalent to the condition that $\mathbb{P}_{0}(\overline{X}^{(1)}_{1}<\infty)=1$. By Borell-TIS inequality, see [3, Theorem 2.1.1], for any $x>0$
(2.1) $\mathbb{P}_{0}\left(\left|\overline{X}^{(1)}_{t}-m(t)\right|>x\right)$ $\leqslant 2e^{-x^{2}/2\sigma(t)^{2}},$
@@
$\mathrm{Var}\left[\overline{X}^{(1)}_{t}\right]\leqslant\int_{0}^{\infty}4xe^{-x^{2}/2\sigma(t)^{2}}=4\sigma(t)^{2}.$
Therefore, $\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{2}\right]\leqslant m(t)^{2}+4\sigma(t)^{2}$. From [10, Equation (29)] it is known that $\lim_{t\downarrow 0}{\sigma(t)^{2}}/{m(t)}=0$. Applying Proposition 1.3, we obtain the following result.
@@
If $\Omega$ is a bounded open set with $C^{1,1}$ boundary then,
(2.2) $\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega).$
@@
In [10, Theorem 4.4] under the condition that the rescaled process $Y^{(t)}_{u}=\mu^{-1}_{t}X^{(1)}_{tu}$ converges weakly to a Gaussian process $(Y_{u})_{u\geqslant 0}$ as $t\downarrow 0$ in Skorokhod $J_{1}$ topology, it is proved that
$\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathbb{E}\left[\overline{Y}_{\!1}\right]\mathrm{Per}(\Omega),$
with $\mathbb{E}[\overline{Y}_{\!1}]\leqslant 1$. Proposition 2.1 does not require the convergence of the rescaled Gaussian process, and hence it strengthens [10, Theorem 4.4].
2.3. Time-changed $\alpha$-stable process
Consider the $d$-dimensional $\alpha$-stable isotropic Lévy process $Y=(Y_{t})_{t\geqslant 0}$ with $\alpha\in(1,2]$, and a nonnegative, increasing process $U=(U_{t})_{t\geqslant 0}$ with càdlàg paths which is independent of $Y$ and $U_{0}=0$. Define $X_{t}=Y_{U_{t}}$. Then, $X=(X_{t})_{t\geqslant 0}$ is translation invariant and isotropic. Let us assume that for some $p>1$,
(2.3) $\mathbb{E}[U^{p/\alpha}_{t}]<\infty\quad\text{for all $t\geqslant 0$},\quad\lim_{t\downarrow 0}\frac{\mathbb{E}[U^{p/\alpha}_{t}]}{\mathbb{E}[U^{1/\alpha}_{t}]}=0.$